A pipe whose average roughness is 0.02mm is employed to raise water through a height of 10m. The pipe length is 100m and its diameter is 300mm. Calculate the power required for a discharge rate of 30,000 l/min
1
Expert's answer
2021-09-19T00:20:26-0400

CQ =0.64, CP =0.60, CH =0.75

Q =CQnD3 +0.64 ·30 s−1 ·(0.5 m)3 =2.40 m3/s

∆h = CH n2 D2

g = 0.75 ·(30 s−1)2 ·(0.5 m)2

9.81 m/s2 =17.2 m.

P =CPρD5n3 =0.60 ·1000 kg/m3 ·(0.5 m)5 ·(30 s−1)3 =506 kW.